Please kindly help me with these 2 questions.
1. Leon drove 270 miles to the lodge in the same time as Pat drove 330 miles to the lodge. If Pat drove 10 miles per hour faster than Leon, then how fast did each of them drive?
2. The student-teacher ratio for Washington High was reported to be 27.5 to 1. If there are 42 teachers, then how many studentd are here?
Thank you

Answer by checkley71(8403)

330/(R+10)=270/R CROSS MULTIPLY
270R+2700=330R
270R-330R=-2700
-60R-2700
R=-2700/-60
R=45 MPH FOR LEON'S SPEED
45+10=55 MPH FOR PAT'S SPEED.
PROOF
330/55=270/45
6=6